OpenStudy (anonymous):

need help with 2 questions differential equations: 1) y' = (2x-y^2) / (y-2xy) 2) (x^3)y' = 2(y^2 + (x^2)y - x^4)

OpenStudy (anonymous):

Try using the app mathway, it's a good calculator for problems like this and is super easy to use :)

OpenStudy (anonymous):

i checked it out, but that doesn't even do differential equations....

OpenStudy (anonymous):

Have you solved exact equations before?

OpenStudy (anonymous):

yes, i tried the method but My and Nx are not equal to each other or am i wrong?

OpenStudy (anonymous):

Have you tried finding an integrating factor to make them exact?

OpenStudy (anonymous):

i tried, but i became too complex

OpenStudy (anonymous):

hello, i believe i got number 1, any help in number two would be appreciated.

OpenStudy (anonymous):

\[\begin{align*}x^3y'&=2(y^2+x^2y-x^4)\\\\ \frac{1}{y^2}y'&=\frac{2}{x^3}+\frac{2}{xy}-2x \end{align*}\] Let \(t=\dfrac{1}{y}\), so that \(t'=-\dfrac{1}{y^2}y'\). \[\begin{align*}-t'&=\frac{2}{x^3}+\frac{2}{x}t-2x\\\\ t'-\frac{2}{x}t&=2x-\frac{2}{x^3} \end{align*}\] which is linear in \(t\).

OpenStudy (anonymous):

As for the first problem, using the exact equation method, \[\begin{align*} y'&=\frac{2x-y^2}{y-2xy}\\\\ y^2-2x+(y-2xy)y'&=0 \end{align*}\] Partial derivatives: \[M_y=2y\\ N_x=-2y\] Integrating factor: \[\begin{align*}\ln\mu(x)&=\int\frac{M_y-N_x}{N}\,dx\\\\&=\int\frac{2y-(-2y)}{y-2xy}\,dx\\\\&=\int\frac{4}{1-2x}\,dx\\\\ &=-2\ln(1-2x)\\\\ \mu(x)&=\frac{1}{(1-2x)^2} \end{align*}\] So the ODE is now \[\frac{1}{(1-2x)^2}(y^2-2x)+\frac{1}{(1-2x)^2}(y-2xy)y'=0\] with partial derivatives, \[{M^*}_y=\frac{2y}{(1-2x)^2}\\ {N^*}_x=\frac{2y}{(1-2x)^2}\]

OpenStudy (anonymous):

thanks a lot!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

A simpler way for solving the first equation involves a substitution, \(y=\sqrt z\), which gives \(y^2=z\) and \(y'=\dfrac{z'}{2\sqrt z}\). \[\begin{align*} y'&=\frac{2x-y^2}{y-2xy}\\\\ \frac{1}{2\sqrt z}z'&=\frac{2x-z}{\sqrt z(1-2x)}\\\\ \frac{1}{2}z'&=\frac{2x-z}{1-2x}\\\\ z'+\frac{2}{1-2x}z&=\frac{4x}{1-2x} \end{align*}\] which is linear in \(z\).

OpenStudy (anonymous):

do you mind looking at this, i used the exact method, i also used symbolab calculator for partial derivatives and integrals, Im not sure if its correct.

OpenStudy (anonymous):

first picture might be small but you can ctrl + +, to zoom in

OpenStudy (anonymous):

Also your second method is not correct, you second to last step to last step is not possible.

OpenStudy (anonymous):

nvm its correct lol

OpenStudy (anonymous):

In your work, your partial derivatives should match up. As above, you have \[{M^*}=\frac{y^2-2x}{(1-2x)^2}~~\implies~~{M^*}_y=\frac{2y}{(1-2x)^2}\\ N^*=\frac{y-2xy}{(1-2x)^2}=\frac{y}{1-2x}~~\implies~~{N^*}_x=\frac{2y}{(1-2x)^2}\] You're looking for a solution of the form \(\Psi(x,y)=C\), where \(\Psi_x=M^*\) and \(\Psi_y=N^*\dfrac{dy}{dx}\). Integrating, you have \[\Psi=\int N^*\,dy=\int\frac{y}{1-2x}\,dy=\frac{y^2}{2(1-2x)}+f(x)\] Differentiating with respect to \(x\) yields \[\Psi_x=M^*=\frac{y^2-2x}{(1-2x)^2}=\frac{y^2}{(1-2x)^2}+f'(x)\] which admits the ODE \[f'(x)=-\frac{2x}{(1-2x)^2}\] Integrating yields \[f(x)=-2\int\frac{x}{(1-2x)^2}\,dx=-\frac{1}{2(1-2x)}-\frac{1}{2}\ln(1-2x)+C\] and so the final solution should be \[\Psi(x,y)=\frac{y^2}{2(1-2x)}-\frac{1}{2(1-2x)}-\frac{1}{2}\ln(1-2x)=C\]

OpenStudy (anonymous):

This can of course be rewritten as needed. \[\Psi(x,y)=\frac{y^2-1}{2(1-2x)}-\ln\sqrt{1-2x}=C\] etc.

OpenStudy (anonymous):

i already handed in the hw, but letting you know, the second to last step is incorrect, under integrating yields. integration and distribution is not correct.

OpenStudy (anonymous):

anyways, ty very much, you are good with this.

OpenStudy (anonymous):

Oh sorry, I must have missed that factor of \(x\) in the numerator...